graph the equation by plotting points
y=x-5

Answer by ewatrrr(24785)

 
Hi,
graph the equation by plotting points
y=x-5
y = 0 -5
0 = x -5
Pt(0,-5) and Pt(5,0) on this line
Plotting points and connecting with a Line